# Okapi: Efficiently Safeguarding Speculative Data Accesses in Sandboxed Environments

### v2

@@ 3 Security Objective
We assume a program $p$ is vanilla-sandboxed, i.e., the high memory locations do not influence the architectural state under sequential program semantics. The program $p$ is breakout-secure or generally-sandboxed if it also satisfies weak speculative non-interference. This means that the high locations do not influence the architectural state under speculative semantics [24]. Provided that high and low locations do not share memory pages (i.e., vanilla sandboxes are implemented page-aligned), Okapi is guaranteed to enforce breakout security for $p$, as described in Sec. 4-A.
@@
In contrast to breakout attacks, poisoning attacks target a victim program $p$ that has access to high memory locations. Additionally, $p$ has measures in place to protect high memory locations from being leaked sequentially. Nonetheless, an attacker can poison the control flow in $p$ to circumvent these measures under speculative semantics and leak confidential data [12, 34].
The victim program can be partitioned into two sets of sections $p_{H}$ and $p_{L}$. All sections of $p$ accessing high memory locations sequentially belong to $p_{H}$, while $p_{L}$ is the complement of $p_{H}$. In other words, ${p_{L}}$ is vanilla-sandboxed w.r.t. the high memory locations of $p$. The program must be partitioned in a way such that each individual section of $p_{H}$ cannot leak more than the data it accesses sequentially. Under such a partitioning, $p$ is poisoning-secure iff ${p_{L}}$ is breakout-secure w.r.t. the high memory locations. A trivial partitioning selects $p_{H}$ to contain only all load instructions to high locations. Sec. 4-B addresses the measures Okapi provides for poisoning security and discusses how to find an efficient and effective partitioning of $p$ into $p_{H}$ and $p_{L}$.
@@
Guarantees for speculative non-interference [24], i.e., ensuring that transient execution side channels do not reveal data in the register file and instruction operands (data in transit), are outside the scope of this paper. In addition, we exclude classical side channel attacks, e.g., monitoring the instruction cache footprint of square-and-multiply exponentiation in RSA [47]. Moreover, physical side channels, including power or electromagnetic side channels, fall outside the scope of this paper.
@@
This security objective covers all TES attacks initialized by speculative loads, independent of the microarchitectural side channel being used to transmit and recover secrets (e.g., cache side channels [28], port contention [4], or data-dependent timing in functional units [3]). In particular, it covers the formidable Spectre-PHT and Spectre-BTB attacks for which, to date, no satisfactory mitigations exist in industrial practice.
4 Okapi Concept
@@
Common to all Spectre variants targeting data at rest is a speculative load instruction that accesses a memory location that could not be accessed by the sequential program flow in that section of code. Preventing the execution of this transient access instruction is sufficient to avert an attack. Blocking all speculative loads from executing would achieve this goal but comes with a significant performance penalty, as shown in [27]. Most of the load instructions executed speculatively are not part of an attack. For minimum performance penalty, it is therefore crucial to only block the malicious ones.
@@
Okapi builds upon this observation by blocking as few speculative load instructions as possible, taking into account legal accesses by the current trust domain. Our proposed architecture leverages the spatial and temporal locality in memory accesses, which means that most of the data retrieved within a short period of time is located in a narrow region in the memory space. Fig. 2 shows, for the SPEC CPU2017 rate suite, that programs indeed fulfill this locality assumption as they access only about 20 pages between context switches, on average. In fact, all benchmarks access at most 60 pages per context switch except for the first context switch, during which around 300 pages are accessed to load the benchmark binaries into memory.
@@
In contrast to other isolation mechanisms, such as in-process isolation [43, 38], Okapi is not aware of a program’s address space and sandbox boundaries in advance and learns about legal accesses by tracking committed load instructions during the execution. Thus, a trust domain denotes the (dynamically growing) set of pages that the currently running thread has, so far, accessed legally and that can therefore also be accessed speculatively.
270270270270300300300300330330330330Total Code PagesCode Pages Mean per Context SwitchTotal Data PagesData Pages Mean per Context SwitchAverageperlbench_0perlbench_1gcc_0gcc_1gcc_2bwaves_0bwaves_1mcfcactuBSSNlbmxalancbmkdeepsjengimagickleelanabexchange2fotonik3droms00101010102020202030303030
@@
Figure 2: The average number of accessed pages (code and data) in between context switches and the total number of allocated pages for the SPEC CPU2017 rate suite.
4-A Achieving Breakout Security
@@
Okapi allows speculative access only to memory pages that have already once been accessed non-speculatively by the same sandbox. This confines potential transient attacks to only pages belonging to the current trust domain and removes all universal read gadgets capable of speculatively leaking data at arbitrary memory addresses by breaking out of the trust domain.
@@
Okapi enforces speculative read permissions with a single safe access bit added to each page table entry in the data TLB. Whenever a page is accessed non-speculatively, the bit is set. Speculative loads are only allowed to obtain their physical address from a TLB entry if the corresponding safe access bit is set. Otherwise, the address translation is blocked until the load becomes non-speculative. This measure effectively prevents Spectre gadgets, such as the ones shown in Fig. 1, from accessing memory locations that do not have the safe access bit set and prevents breakout attacks for page-aligned sandboxes. In many cases, the first access to a page causes a miss in the TLB and subsequently triggers an expensive page table walk. Therefore, the delay induced by this mechanism in Okapi is largely amortized by the page table walk which reduces the overall performance overhead incurred.
@@
During execution, the program architecturally accesses data pages and enables the corresponding safe access bits. Therefore, the trust domain, i.e., the amount of data that can possibly be leaked transiently, grows over time. Resetting the safe access bits at the right point in time is crucial to prevent the (mis-)use of stale access rights. The size of a program that establishes a trust domain can vary significantly, ranging from single functions to entire threads or processes. For correct handling of the access bits in Okapi, transitions between programs must be transparent to the CPU.
Sandboxes at process or thread level can easily be implemented in hardware. Privilege switches (cf. Sec. 2-B) are detected by monitoring the corresponding Control and Status Registers (CSRs) and automatically clear the safe access bits.
Therefore, as a first important result, we note that the Okapi hardware alone inhibits all universal read gadgets (e.g., the ones exploited by Spectre-PHT, -BTB and -STL, as depicted in Fig. 1) and achieves breakout security for thread-sized sandboxes.
In practice, it can be desirable to have more than one sandbox within the same thread or process to save costly context switches [15]. Detecting transitions from one sandbox to another below thread granularity is, unfortunately, not immediately possible at the hardware level. On the other hand, at the software level, sandboxes of any size are clearly distinguished. Therefore, Okapi empowers the software to communicate sandbox exits to the hardware.
@@
We add a dedicated instruction, OkapiReset, to the ISA that resets the trust domain by flushing all safe access bits. The OkapiReset instruction acts as a speculation barrier for load instructions. Similar to other speculation barriers, such as fences, it is executed only upon reaching the head of the Re-Order Buffer (ROB) to avoid premature resetting. As long as an OkapiReset instruction is present in the ROB, all younger load instructions are serialized to prevent out-of-order use of existing access permissions. Software developers can make use of the OkapiReset instruction to establish more fine-grained sandboxes, e.g., at function level.
@@
4-B Achieving Poisoning Security
While establishing trust domains for programs of a chosen granularity prevents transient accesses to the outside, data located on pages with an activated safe access bit can still be loaded speculatively. The challenge for Okapi w.r.t. such poisoning attacks (cf. Sec. 3) is to guarantee that those code sections $p_{L}$ within a program that do not sequentially access confidential data cannot misuse speculative access permissions gained by sensitive code sections $p_{H}$. We need to ensure that upon a transition from $p_{H}$ to $p_{L}$ the high memory locations are not contained in the trust domain. Placing confidential and public data on separate pages and augmenting the program with an OkapiReset instruction after every access to confidential data solves this problem as it effectively constrains $p_{H}$ to only contain all confidential load instructions. However, this incurs a substantial performance overhead as every OkapiReset instruction revokes all safe access bits, thereby also hindering speculation on previously accessed public data.
We therefore propose a second Okapi instruction which can preserve speculation on previously accessed public data. We introduce the OkapiLoad instruction that allows the software to access confidential data pages non-speculatively without adding the page to the trust domain, i.e., not setting the safe access bit. This ensures poisoning security as defined in Sec. 3 if every access to confidential data is performed by an OkapiLoad instruction.
Achieving full poisoning security for sandboxes that repeatedly process confidential data with these instructions comes at a high overhead due to serializing accesses to high memory locations. This can be relaxed by increasing the size of the sections in $p_{H}$ from single load instructions to larger segments of code to be able to speculate on high memory locations without leaking them speculatively.
A major challenge for increasing the size of the sections in $p_{H}$ are Spectre-BTB attacks, illustrated in Fig. 3. By poisoning branch prediction, an attacker can transiently divert the control flow from the sequential program flow (green arrow) to a gadget (blue arrow). Without additional measures, transiently executed gadgets can still access (red arrow) the content of pages already marked with a safe access bit, i.e., data pages accessed by $p_{H}$. This gadget can reside anywhere in the mapped code space, posing a challenge to conventional mitigations [46].
@@ 4-B Achieving Poisoning Security
We can enlarge the sections in $p_{H}$ as long as they fulfill the property of not containing any gadgets (cf. Sec. 3). Therefore, if we can guarantee a candidate region to be free of gadgets [46], we can extend the sections in $p_{H}$ to a chosen granularity, e.g., an entire function. The inspection of candidate regions can be done individually.
In conclusion, Okapi offers two different strategies to block poisoning attacks in sandboxes. The developer can either prevent confidential data pages from entering the trust domain by using the OkapiLoad instruction or revoke the access rights via an OkapiReset instruction when switching from $p_{H}$ to $p_{L}$. Note that already the hardware measures by themselves drastically reduce the code space available for exploitable gadgets as shown in our experiments in Sec. 9-B.
@@
Trust Domain <Data Page> <Data Page> <Data Page> ... Secret data ... ... ... ... <Code Page> ... Jump to Func ... ... Func ... <Code Page> ... ... ... SpectreGadget ... ... poisoned branch!
@@
Figure 3: Spectre-BTB challenge for Okapi. If speculative execution flow leaves the current page, all loads on the suspicious page are delayed until accesses to that code page are no longer speculative.
5 Security Reasoning
@@
Okapi prevents speculative leakage of data outside a program’s memory space. The program counter of committed instructions as well as data and address of committed loads and stores of the program are considered exposed and therefore not protected. Compared to schemes that enforce weak speculative non-interference at an address granularity [48, 37, 45], Okapi achieves significantly better performance by declassifying an entire data page with the first non-speculative access to this page. As long as no OkapiLoad instruction is used, the Okapi hardware dynamically labels sequentially accessed data pages as low, i.e., public. This allows for a highly efficient protection against cross-sandbox breakout attacks solely in hardware.
@@
As a result of the native hardware support for virtual memory translation based on page tables (cf. Sec. 2-C), most operating systems and runtime environments, such as in-process isolation [38, 35] and enclaves [7], utilize page granularity to define the accessible address space and enforce software boundaries. Commercial cloud services, such as Cloudflare workers [14], similarly incorporate page-aligned partitioning of the address space between mutually distrusting tenants. Therefore, mutually distrusting software domains typically possess non-overlapping page-aligned data memory address spaces, and page sharing only happens in instruction memory through shared libraries. Therefore, in the context of sandboxing, secure speculation with finer than page granularity has limited benefits while incurring a substantial performance penalty.
@@
”Classic” Sandboxing
In classic sandbox scenarios where a trusted environment is protected from being accessed by an untrusted instance, we only need to ensure that all safe access bits are reset when handing over control to the untrusted job. If switching to the untrusted job involves a privilege switch, this is already taken care of by the Okapi hardware. Otherwise, an OkapiReset instruction needs to revoke all previously gained access rights.
@@
Securing runtime environments that run multiple tenants within a single thread against breakout attacks is a major concern in industrial applications, e.g., serverless computing [40, 14]. Okapi secures this kind of sandboxed environments by preventing speculative cross-domain accesses even within the same virtual address space. Therefore, the runtime environment, e.g., the WebAssembly engine, needs only minor modifications to execute an OkapiReset instruction each time it enters and exits a sandbox, to fully separate the runtime and each sandbox from each other.
@@
The inverse scenario to ”classic” sandboxing is vaulting, where a program is protected against malicious accesses from the outside. Besides breaking into the vault, attackers can launch a poisoning attack (cf. Sec. 3) to trick the victim program into speculatively leaking its own confidential data. Okapi already covers breakouts from other programs into vaulted code segments with the security mechanisms protecting mutually distrusting sandbox environments. In addition, Okapi mitigates poisoning attacks by employing the hardware measures described in Sec. 4-B. If desired, the OkapiLoad instruction prevents confidential data from being added to the vault’s trust domain.
6 Okapi Architecture
6-A Hardware Architecture
Trust DomainFetchDecodeRenameExecuteWritebackCommit1273ROB IQ 6 LSQ 45TLBsafe access bitInstruction Cache PortData Cache Port
Figure 4: Out-of-Order processor augmented with Okapi.

### v3

@@ 3. Security Objective
We assume that a program $p$ is sandboxed in software (vanilla sandboxed), i. e., the high memory locations do not influence the architectural state under  $p$’s sequential program semantics. The program $p$ is breakout-secure or generally sandboxed if it also satisfies weak speculative non-interference. This means that the high locations do not influence the architectural state even under speculative semantics (Guarnieri et al., 2021). Provided that high and low locations do not share memory pages (i. e., software sandboxes are implemented page-aligned), Okapi is guaranteed to enforce breakout security for $p$, as described in Sec. 4.1.
@@
In contrast to breakout attacks, poisoning attacks target a victim program $p$ that has access to high memory locations. Additionally, $p$ has measures in place to protect high memory locations from being leaked sequentially, e.g., by relying on constant-time code for sensitive computations. Nonetheless, an attacker can poison the control flow in $p$ to circumvent these measures under speculative semantics and leak confidential data (Cauligi et al., 2022; Narayan et al., 2021).
The victim program can be partitioned into two sets of sections, $p_{H}$ and $p_{L}$. All sections of $p$ accessing high memory locations sequentially belong to $p_{H}$, while $p_{L}$ is the complement of $p_{H}$. In other words, $p_{L}$} never accesses confidential data and is therefore vanilla sandboxed (Guarnieri et al., 2021) w. r. t. the high memory locations of $p$. The program must be partitioned in a way such that each individual section of $p_{H}$ cannot leak more than the data it accesses sequentially. Under such a partitioning, $p$ is poisoning-secure iff $p_{L}$} is breakout-secure w. r. t. the high memory locations. A trivial partitioning selects $p_{H}$ to contain only all load instructions to high locations. Sec. 4.2 addresses the measures Okapi provides for poisoning security and discusses how to find an efficient and effective partitioning of $p$.
@@
Guarantees for speculative non-interference (Guarnieri et al., 2021), i. e., ensuring that transient execution side channels do not reveal data in the register file and instruction operands (data in transit), are outside the scope of this paper. This includes side channels due to predictive store forwarding (AMD, 2023) and other optimizations that can leak the content of the store buffer. In addition, we exclude classical side channel attacks, e. g., monitoring the instruction cache footprint of square-and-multiply exponentiation in RSA (Yarom and Falkner, 2014) or data-dependent prefetcher-based attacks violating constant-time guarantees (Chen et al., 2024).
Moreover, physical side channels, including power or electromagnetic side channels, fall outside the scope of this paper.
@@
Our security objective covers all TES attacks initialized by speculative loads, independent of the microarchitectural side channel being used to transmit and recover secrets (e. g., cache side channels (Kocher et al., 2019), port contention (Behnia et al., 2021), or data-dependent timing in functional units (Andrysco et al., 2015)). In particular, it covers the formidable Spectre-PHT and Spectre-BTB attacks for which, to date, no satisfactory mitigations exist in industrial practice.
4. Okapi Concept
@@
Common to all Spectre variants targeting data at rest is a speculative load instruction that accesses a memory location that could not be accessed by the sequential program flow. Preventing the execution of this transient access instruction is sufficient to avert an attack. Blocking all speculative loads from executing would achieve this goal but comes with a significant performance penalty, as shown in (Jauch et al., 2023). Most of the load instructions executed speculatively are not part of an attack. For minimum performance penalty, it is therefore crucial to only block the malicious ones.
@@
Okapi builds upon this observation by blocking as few speculative load instructions as possible, taking into account legal accesses by the current trust domain. Our proposed architecture leverages the spatial and temporal locality in memory accesses, which means that most of the data retrieved within a short period of time is located in a narrow region in the memory space. Fig. 2 shows, for the SPEC CPU2017 rate suite, that programs indeed fulfill this locality assumption as they access only about 20 pages between context switches, on average. In fact, all benchmarks access at most 60 pages per context switch except for the first context switch, during which around 300 pages are accessed to load the binaries into memory.
@@
In contrast to other isolation mechanisms, such as in-process isolation (Vahldiek-Oberwagner et al., 2019; Schrammel et al., 2020), Okapi is not aware of a program’s address space and sandbox boundaries. Instead, it supplies a secure hardware basis that guarantees the enforcement of boundaries provided by security mechanisms in software. The Okapi hardware learns about legal accesses by tracking committed load instructions during the execution. Thus, an Okapi trust domain comprises the (dynamically growing) set of pages that the currently running thread has, so far, accessed legally and that it may, therefore, also access speculatively.
[…]
The figure shows a bar plot. On the x-axis, most of the SPEC CPU2017 benchmark programs are listed. The y-axis displays the number of accessed memory pages these programmes accessed. Per program, four bars are given: the total number of code pages, the total number of data pages, the code pages mean per context switch, and the data pages mean per context switch. On average, each programm access a total number of around 280 code pages and around 24 data pages. Per context switch, however, the each program only accesses around 18 code pages and around 8 data pages.
@@
Figure 2. The average number of accessed pages (code and data) in between context switches and the total number of allocated pages for the SPEC CPU2017 rate suite.
4.1. Achieving Breakout Security
@@
Okapi allows speculative access only to memory pages that have already been accessed non-speculatively at least once by the same sandbox. This confines potential transient attacks to only the pages belonging to the current trust domain and removes all universal read gadgets capable of speculatively leaking data at arbitrary memory addresses by breaking out of the trust domain.
@@
Okapi enforces speculative read permissions with a single safe access bit added to each page table entry in the data TLB. Whenever a page is accessed non-speculatively, the bit is set. Speculative loads are only allowed to obtain their physical address from a TLB entry if the corresponding safe access bit is set. Otherwise, the address translation is blocked until the load becomes non-speculative. This measure effectively prevents Spectre gadgets from accessing memory locations that do not have the safe access bit set and thwarts breakout attacks for page-aligned sandboxes. In many cases, the first access to a page causes a miss in the TLB and triggers an expensive page table walk. Hence, the delay induced by this mechanism in Okapi is largely shadowed by the page table walk which reduces the overall performance overhead incurred by Okapi.
In this paper, we describe the Okapi mechanism for a single core without multi-threading, featuring a physically indexed, physically tagged cache. For other system configurations, including ones with simultaneous multi-threading (SMT), the functionality can be adapted accordingly. For example, a separate safe access bit can be implemented per SMT thread.
Note that data-dependent prefetchers need to perform an address translation for dereferencing potential pointers (Vicarte et al., 2022), i.e., the prefetcher requests can be blocked with the same mechanism in the TLB. Thus, data-dependent prefetchers do not pose an additional breakout threat to data at rest (cf. Sec. 3).
@@
During execution, the program successively accesses data pages architecturally and sets the corresponding safe access bits. Accordingly, the trust domain, i. e., the number of memory pages from which data can possibly be leaked transiently, grows over time. Resetting the safe access bits at the right point in time is crucial to prevent the (mis-)use of stale access rights. The size of a sandbox that establishes a trust domain can vary significantly, ranging from single functions to entire threads or processes. For correct handling of the access bits in Okapi, transitions between trust domains must be transparent to the CPU.
Sandboxes at process or thread level can easily be implemented in hardware. Privilege switches (cf. Sec. 2.2) are detected by monitoring the corresponding Control and Status Registers (CSRs) and automatically clear the safe access bits.
Therefore, as a first important result, we note that the Okapi hardware alone inhibits all universal read gadgets (e. g., the ones exploited by Spectre-PHT, -BTB and -STL, as depicted in Fig. 1(c)) and achieves breakout security for thread-sized sandboxes.
In theory, to prevent breakout attacks one could create a new process for every sandbox. In practice however, it is desirable to have more than one sandbox within the same thread or process to avoid the high cost of context switches (Cloudflare, 2024).
When multiple sandboxes are mapped to the same process, the straightforward approach of delaying load instructions in case of a TLB miss (Sakalis et al., 2019) is not sufficient to guarantee secure speculation. The first sandbox populates the TLB with its entries and subsequent sandboxes can re-use these entries to perform a breakout attack.
While sandboxes are clearly distinguishable in the software, detecting transitions between sandboxes below thread level is not directly possible in the hardware. For this reason, Okapi introduces a special instruction using which the software can communicate sandbox exits to the hardware.
@@
We add a dedicated instruction, OkapiReset, to the ISA that resets the trust domain by flushing all safe access bits. The OkapiReset instruction acts as a speculation barrier for load instructions. Similar to other speculation barriers, such as fences, it is executed only upon reaching the head of the Re-Order Buffer (ROB) to avoid premature resetting. As long as an OkapiReset instruction is present in the ROB, all younger load instructions are serialized to prevent out-of-order use of existing access permissions. Software developers can make use of the OkapiReset instruction to implement more fine-grained sandboxes, e. g., at function level.
@@
4.2. Achieving Poisoning Security
While establishing trust domains for programs of a chosen granularity prevents transient accesses to the outside, data located on pages with an activated safe access bit can still be loaded speculatively. The challenge for Okapi w. r. t. poisoning attacks (cf. Sec. 3) is to guarantee that code sections $p_{L}$ within a program that do not sequentially access confidential data cannot misuse speculative access permissions gained by sensitive code sections $p_{H}$. We need to ensure that upon a transition from $p_{H}$ to $p_{L}$ the high memory locations are not contained in the trust domain. Placing confidential and public data on separate pages and augmenting the program with an OkapiReset instruction after every access to confidential data solves this problem as it effectively constrains $p_{H}$ to contain all confidential load instructions. However, this incurs a substantial performance overhead as every OkapiReset instruction revokes all safe access bits, thereby also preventing speculation on previously accessed public data.
We therefore introduce the OkapiLoad instruction that allows the software to access confidential data pages non-speculatively without adding the page to the trust domain, i. e., not setting the safe access bit. This ensures poisoning security (cf. Sec. 3) if every access to confidential data is performed by an OkapiLoad and preserves speculative access rights on previously accessed public data.
Achieving full poisoning security for sandboxes that repeatedly process confidential data with these instructions comes at a high overhead due to serializing accesses to high memory locations. This can be relaxed by increasing the size of the sections in $p_{H}$ from single load instructions to larger segments of code to be able to speculate on high memory locations without leaking them.
A major challenge for increasing the size of the sections in $p_{H}$ are Spectre-BTB attacks, illustrated in Fig. 3. By poisoning branch prediction, an attacker can transiently divert the control flow from the sequential program flow (green arrow) to a gadget (blue arrow). Without additional measures, transiently executed gadgets can still access (red arrow) the content of pages already marked with a safe access bit, i. e., data pages accessed by $p_{H}$. This gadget can reside anywhere in the mapped code space, posing a challenge to conventional mitigations (Wiebing et al., 2024).
@@ 4.2. Achieving Poisoning Security
We can enlarge the sections in $p_{H}$ as long as they fulfill the property of not containing any gadgets (cf. Sec. 3). Therefore, if we can guarantee a candidate region to be free of gadgets (Wiebing et al., 2024), we can extend the sections in $p_{H}$ to a chosen granularity, e. g., an entire function. The inspection of candidate regions can be done independently of each other.
In conclusion, Okapi offers two different strategies to blocking poisoning attacks in sandboxes. The developer can either prevent confidential data pages from entering the trust domain by using the OkapiLoad instruction or revoke the access rights via an OkapiReset instruction when switching from $p_{H}$ to $p_{L}$.
Note that the hardware measures alone drastically reduce the code space available for exploitable gadgets as shown in our experiments in Sec. 9.2.
@@
4.3. Okapi’s Security Guarantees
In summary, the features described in Sec. 4.1 and Sec. 4.2 enable Okapi to effectively and efficiently protect against breakout and poisoning attacks. In particular, Okapi eliminates all threats imposed by Spectre-like universal read gadgets with only minor performance and hardware penalty (cf. Sec. 7 & Sec. 8). Note that this security guarantee is independent of the classification of different Spectre variants. Okapi prevents unsafe access instructions which are a necessary component of every Spectre attack, independent of the covert channel being used. Hence, it covers all Spectre variants targeting data residing in memory.
In addition, Okapi by design eliminates Meltdown- and MDS-style vulnerabilities since it permits access to physical memory addresses only in case of a non-faulting virtual address translation (cf. Sec. 9). Thus, attackers cannot bypass any kind of memory exception which is the underlying mechanism for these attacks.
The hardware features incur low performance and implementation overhead and provide breakout security for thread-sized sandboxes (cf. Sec. 9.1). In addition, without further measures, they already drastically limit the set of possible locations for gadgets exploitable in poisoning attacks.
With additional security measures at the software level leveraging the OkapiReset instruction, Okapi delivers breakout security in custom-sized sandboxes (cf. Sec. 7.2). Full security against poisoning attacks can be achieved by identifying and hardening gadgets solely within the limited code region defined by the victim sandbox’s code. Alternatively, access to secret data can be controlled using the OkapiLoad instruction. This, at the cost of additional performance overhead, relieves the software developer from manually identifying gadgets within the victim’s code base (cf. Sec. 7.3).
Both Okapi instructions act like pseudo-fences. They serialize younger load instructions and thus do not cause any additional data-dependent timing. Furthermore, load instructions in Okapi only access addresses on non-faulting pages and do therefore not create any secret-dependent behavior. As a result, Okapi does not compromise any existing constant-time guarantees.
[…]
This figure shows three data pages that currently belong to the trust domain and contain secret data. Also, two code pages are show: one contains the code that is currently being executed and contains an indirect jump, and one page that contains a Spectre gadget. Arrows indicate Okapis behavior when an attacker poisons the branch predictor to divert the indirect jump to speculatively execute the Spectre gadget on the second code page. Since the gadget lies on a different page, every load performed by the gadget gets flagged as suspicious and gets withheld until its execution becomes save. This does not happen since the gadget does not execute architecturally and therefore Okapi effectively blocks the gadget’s access to the secret.
@@ 4.3. Okapi’s Security Guarantees
Figure 3. Spectre-BTB challenge for Okapi. If speculative execution flow leaves the current page, all loads on the suspicious page are delayed until accesses to that code page are no longer speculative.
5. Okapi in different threat scenarios
@@
Okapi prevents speculative leakage of data outside a program’s memory space. Compared to schemes that enforce weak speculative non-interference at an address granularity (Yu et al., 2019; Sakalis et al., 2019; Weisse et al., 2019), Okapi achieves significantly better performance by declassifying an entire data page with the first non-speculative access to this page. As long as no OkapiLoad instruction is used, the Okapi hardware dynamically labels sequentially accessed data pages as low, i. e., public. This allows for a highly efficient protection against cross-sandbox breakout attacks solely in hardware.
@@
As a result of the native hardware support for virtual memory translation based on page tables (cf. Sec. 2.3), most operating systems and runtime environments, such as in-process isolation (Schrammel et al., 2020; Narayan et al., 2023) and enclaves (Bourgeat et al., 2019), utilize page granularity to define the accessible address space and enforce software boundaries. Commercial cloud services, such as Cloudflare workers (Cloudflare, 2018), similarly incorporate page-aligned partitioning of the address space between mutually distrusting tenants. Therefore, mutually distrusting software domains typically possess non-overlapping page-aligned data memory address spaces, and page sharing only happens in instruction memory through shared libraries. Thus, in the context of sandboxing, secure speculation with finer than page granularity has limited benefits while incurring a substantial performance penalty.
@@
“Classic” Sandboxing
In classic sandbox scenarios where a trusted environment is protected from being accessed by an untrusted instance, we only need to ensure that all safe access bits are reset when handing over control to the untrusted job. If switching to the untrusted job involves a privilege switch, this is already taken care of by the Okapi hardware. Otherwise, the sandbox runtime must execute an OkapiReset instruction to revoke all previously gained access rights.
@@
Securing runtime environments that run multiple tenants within a single thread against breakout attacks is a major concern in industrial applications, e. g., serverless computing (Schwarzl et al., 2022; Cloudflare, 2018). Okapi secures this kind of environment by preventing speculative cross-domain accesses even within the same virtual address space. Therefore, the runtime environment, e. g., the WebAssembly engine, needs only minor modifications to execute an OkapiReset instruction each time it enters and exits a sandbox, to fully separate the runtime and each sandbox from each other.
@@
The inverse scenario to “classic” sandboxing is vaulting, where a program is protected against malicious accesses from the outside. Besides breaking into the vault, attackers can launch a poisoning attack (cf. Sec. 3) to trick the victim program into speculatively leaking its own confidential data. Okapi already covers breakouts from other programs into vaulted code segments with the security mechanisms protecting mutually distrusting sandbox environments. In addition, Okapi mitigates poisoning attacks by employing the hardware measures described in Sec. 4.2. If desired, the OkapiLoad instruction prevents confidential data from being added to the vault’s trust domain.
Kernel as a Confused Deputy
The Okapi hardware automatically resets the safe access bits on every privilege switch independent from the software. Therefore, confused-deputy attacks against the kernel are automatically limited. As soon as the privilege switch occurs, the kernel can no longer speculatively access data pages belonging to the threads that were running before. Therefore, the kernel can never leak a thread’s information that it does not access architecturally during a privilege switch.
If the attack targets data that is accessed by the kernel sequentially, it can only be successful if there is a gadget on the same instruction page of the kernel code.
6. Okapi Architecture
6.1. Hardware Architecture
[…]
The figure shows a 6 stage CPU pipeline. The stages are: fetch, decode, rename, execute, writeback, and commit. The fetch stage has access to the instruction cache port. After the rename stage, the instruction is forwarded to the execute stage, the re-order buffer (ROB), the instruction queue (IQ), and the load store queue (LSQ), which in turn has access to the translation-lookaside buffer (TLB) with the safe access bits and the data cache port. Some components are marked with numbers that are explained in the text. The fetch stage is marked with one, decode with two, the ROB with three, the LSQ with four, the TLB with five, the IQ with six and the execute stage with seven.
Figure 4. Out-of-Order processor augmented with Okapi.